### v2

@@ 6.1 Mirror Descent
of Theorem 8. According to the descent property in (95), we have $f(\mathbf{x}_{t+1})-f^{*}\leq f(\mathbf{x}_{t})-f^{*}\leq\cdots\leq f(\mathbf{x}_{0})-f^{*}$. By Lemma 18, we deduce that for $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$ holds for all $t\in\mathbb{N}$, where $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2(f(\mathbf{x}_{0})-f^{*})\ell(2\alpha)\right\}$. After establishing such an upper bound for the dual norm of the gradients along the trajectory, we can cope with generalized smooth functions in a similar way as the global smooth functions. Similar to (88), we consider the update of mirror descent in (9):
@@
Note that the local Lipschitz constant $\ell(2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*})$ is bounded by $L=\ell(2G)$. By Lemma 17 and $\eta\leq\frac{1}{L}$, we have
@@
Therefore, the distance between $\mathbf{x}_{t+1}$ and $\mathbf{x}_{t}$ is small enough to satisfy the conditions in Lemma 3. Using the effective smoothness characterization, we obtain
@@
Taking maximum on both sides of the inequality and under Assumption 2, the rate of convergence of the average-iterate is given by:
$f(\overline{\mathbf{x}}_{T})-f^{*}\leq\frac{\max_{\mathbf{x}\in\mathcal{X}}B(\mathbf{x},\mathbf{x}_{0})}{\eta T}\leq\frac{D^{2}}{\eta T}.$ (31)

### v3

@@ Appendix E Analysis for Mirror Descent
According to the descent property in (61), we have $f(\mathbf{x}_{t+1})-f^{*}\leq f(\mathbf{x}_{t})-f^{*}\leq\cdots\leq f(\mathbf{x}_{0})-f^{*}$. By Lemma E.2, we deduce that for $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$ holds for all $t\in\mathbb{N}$, where $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2(f(\mathbf{x}_{0})-f^{*})\ell(2\alpha)\right\}$. After establishing such an upper bound for the dual norm of the gradients along the trajectory, we can cope with generalized smooth functions in a similar way as the global smooth functions. Similar to (54), we consider the update of mirror descent in (6):
@@
Note that the local Lipschitz constant $\ell(2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*})$ is bounded by $L=\ell(2G)$. By Lemma E.1 and $\eta\leq\frac{1}{L}$, we have
@@
Therefore, the distance between $\mathbf{x}_{t+1}$ and $\mathbf{x}_{t}$ is small enough to satisfy the conditions in Lemma 2.7. Using the effective smoothness characterization, we obtain